OpenStudy (anonymous):

tabitha is raising rabbits to sell during the spring. She has 16 rabbits and feeds them each one-half cup of alfalfa every morning. The bag of pellets she bought completely filled a 5 gallon plastic barrel. How many days will it last?

OpenStudy (texaschic101):

16 rabbits at 1/2 cup per day = 16(1/2) = 16/2 = 8 cups per day 1 gal is 16 cups, so 5 gallons is 5(16) = 80 cups so if your feeding the rabbits 8 cups per day, and a 5 gallon barrel holds 80 cups, then the 5 gallon barrel should last 10 days.